# Real-time Monocular Full-body Capture in World Space via Sequential Proxy-to-Motion Learning

### v1

@@ 4.1 Sequential Full-body Motion Recovery
For better body-hand compatibility, we exploit the cross-attention mechanism to facilitate the motion context sharing during the body and hand recovery. Specifically, we first obtain the initial body and hand features from the temporal encoders and split them as the Query $Q_{b},Q_{h}$, Key $K_{b},K_{h}$, and Value $V_{b},V_{h}$. Then, these features will be updated by the cross-attention layer as follows:
$V^{\prime}_{b}$ $=V_{b}+Softmax(\frac{Q_{h}K_{v}^{\top}}{\sqrt{d_{k}}})V_{b},$ (3)
$V^{\prime}_{h}$ $=V_{h}+Softmax(\frac{Q_{b}K_{h}^{\top}}{\sqrt{d_{k}}})V_{h}.$
@@
The updated body and hand features $V^{\prime}_{b},V^{\prime}_{h}$ will then be fed into the corresponding regressors to predict pose parameters. We demonstrate that the fusion strategy can make our regressors benefit each other to produce more comparable wrist poses in the full-body model.
Figure 4: Illustration of the proposed cross-attention for neural motion descent.
@@ 4.2 Contact-aware Neural Motion Descent
At the second stage of the proposed scheme, the coarse motion predictions will be refined to be more accurate and physically plausible with our proposed contact-aware neural motion descent module. This module takes the foot-ground contact and misalignment status as input and produces updating human parameters during iterations.
To refine the predictions, our network needs to perceive the misalignment and the foot-ground contact status of the currently predicted motions $J_{i},v_{i}=\mathcal{M}(\beta_{i},t_{i},\theta_{i})$, where the $J_{i}$ and $v_{i}$ denote the 3D joints and vertices of the body model at the $i$-th iteration. Since we assume the camera is steady, the estimated motion parameters $\beta,t,\theta_{b}$ in Sec. 4.1 should be transformed to the world space with a reference camera pose as the initial predictions $\beta^{0},t^{0},\theta_{b}^{0}$. In practice, we take the predicted camera pose of the first frame as the reference camera pose and then transform subsequent estimated human motions accordingly.
@@
For the misalignment status, we can project the 3D joints on the image plane and calculate the difference between the reprojected 2D joints and the proxy observations: $\mathcal{S}_{proj}=\Pi(R_{c}^{\top}J_{i}+T_{c})-\widehat{J}_{2D}$. For the contact status, we calculate the foot-ground contact based on the vertex velocity $v^{xy}_{i}-v^{xy}_{t-1}$ in XY-plane and the distance $v^{z}_{i}$ to the Z-ground plane. Moreover, we also leverage the temporal features from inputs 2D skeletons to predict the contact labels $ind_{t}$, which will be used as an indicator to mask the foot-ground contact. Then, the contact status of the current predictions can be denoted as $\mathcal{S}_{contact}=ind_{t}\odot(v^{xy}_{i}-v^{xy}_{t-1},v^{z}_{i})$, where $\odot$ denotes the mask operation.
@@
After obtaining the contact and misalignment status, we feed them into the neural motion descent module for motion updates. As shown in Fig. 4, the descent module takes the two groups of tensors as input: i) the state group includes the current SMPL parameters $\beta_{i},t_{i},\theta_{i}$, camera pose $\alpha,\phi,h$ and the sequential motion context $F_{seq}=\{V^{\prime}_{b},V^{\prime}_{h}\}$; ii) the deviation group includes the current misalignment status $\mathcal{S}_{proj}$ and contact status $\mathcal{S}_{contact}$. A straightforward solution would be using an MLP to process these two groups of tensors. However, the values of these two groups exhibit significant differences. For instance, the values of the state tensors change smoothly while the values of the deviation tensors may change rapidly along with the misalignment and contact status. Simply concatenating them as inputs introduce difficulty in the learning process. Note that the magnitude of the deviation tensors is highly correlated with the parameter updates. When the body model is well-aligned without foot skating or ground penetration, the values of the deviation tensors are almost zero, indicating that the refiner should output zeros to prevent further changes in the pose parameters. Otherwise, the refiner should output larger update values for motion adjustments. To leverage such a correlation property, we exploit a cross-attention module to build a more effective architecture. As shown in Fig. 4, two fully-connect layers are leveraged to process the tensors of the state and deviation groups and produce the Query, Key, and Value for the cross-attention module. In this way, our contact-aware neural motion descent module can effectively learn the relationship between the state and deviation groups and hence produce more accurate motion updates. Moreover, the sequential motion context $F_{seq}$ is also leveraged in our neural descent module to mitigate the depth uncertainty and improve the motion predictions.
@@
Compared with previous work [44, 71, 51, 64], the proposed contact-aware neural motion descent module offers the advantage of freeing us from the heavy cost of explicit gradient calculations or the manual tuning of hyper-parameters during testing. Furthermore, the module is capable of learning human motion priors with contact in world space from our synthetic dataset, which provides a more suitable descent direction and step to escape the local minima and achieve faster convergence.
@@ 4.3 Loss Function
In our solution, the full-body motion recovery module and the contact-aware neural motion descent module are trained separately. Benefiting from the proxy-to-motion learning, the ground-truth full-body pose $\theta_{b+h}$ and human body shape $\beta_{b}$ can be obtained for supervision from our synthetic dataset. Overall, the objective function of motion recovery can be written as follows:
$\mathcal{L}_{rec}=\mathcal{L}_{3D}+\mathcal{L}_{2D}+\mathcal{L}_{\theta}+\mathcal{L}_{\beta}+\mathcal{L}_{cam}+\mathcal{L}_{smooth}$ (4)
Specifically, $\mathcal{L}_{3D}$ involves 3D MPJPE loss and 3D trajectory L1 loss while $\mathcal{L}_{2D}$ is the projected 2D MPJPE loss. $\mathcal{L}_{\theta},\mathcal{L}_{\beta},\mathcal{L}_{cam}$ represents L1 loss between the estimated human pose, shape and camera pose to our synthetic ground truth. $\mathcal{L}_{smooth}$ is adopted from  [65] by penalizing the velocity and acceleration between the estimation and the ground truth. For the neural descent module, the objective loss can be written as:
$\begin{cases}\mathcal{L}_{desc}=\sum_{k}{u^{K-k}(\mathcal{L}_{rec}+\mathcal{L}_{contact}+\mathcal{L}_{ind})}\\ \mathcal{L}_{contact}=\sum_{i}{ind^{gt}_{i}\odot(||v^{xy}_{i}||_{2}+||v^{z}_{i}||_{2})}\\ \mathcal{L}_{ind}=\sum_{i}{Entropy(ind^{gt},ind^{est})}\end{cases}$ (5)
where $k=1,2,...,K$ is the iteration time and $u$ is the decay ratio to emphasize the last iteration. We set $K=3,u=0.8$ in practice. $\mathcal{L}_{contact}$ involves the error of trajectory drifting, foot floating, or ground penetration. $\mathcal{L}_{ind}$ refers to the loss between the predicted contact label to the ground truth.
@@ 5 Experiments
In this section, we evaluate the effectiveness of our method on both the synthetic dataset and the public datasets. We also conduct ablation studies to demonstrate the effectiveness of each part and report the computational complexity of each module in our real-time system.
Table 1: Quantitative comparison with state-of-the-art methods on the Human3.6M dataset. *** indicates that the input 2D pose is detected by the cascaded pyramid network (CPN), and $\dagger$ denotes the ground truth 2D pose input.
Method Ground-contact Full-body MPJPE P-MPJPE
Skeleton PoseFormer [74] $\dagger$ 31.3 -
PoseFormer [74] *** 44.3 34.6
VideoPose3D [42] $\dagger$ 37.2 27.2
VideoPose3D [42] *** 46.8 36.5
PoseTriplet [10] *** 78.0 51.8
Ray3D [66] $\dagger$ 34.4 -
Single Frame I2L-MeshNet [34] 55.7 41.1
LearnedGD [51] - 56.4
HUND [64] 69.5 52.6
Pose2Mesh [5] 64.9 46.3
HMR [16] 88.0 56.8
SPIN [21] 62.5 41.1
PyMAF-X [68] 58.1 40.2
Temporal Kanazawa et al. [17] - 56.9
Arnab et al. [1] 77.8 54.3
DSD [53] 59.1 42.4
VIBE [18] 65.9 41.5
TCMR [4] 62.3 41.1
SimPoE [63] 56.7 41.6
DnD [23] 52.5 35.5
Zhou et al. [75] 50.3 -
FrankMocap [46] - -
Ours *** 53.9 42.4
Ours $\dagger$ 40.5 32.3
Figure 5: Visualization of reconstruction results across different cases in the (a) 3DPW [58], (b) EHF [40], and (c) Human3.6M [14] datasets and (d,e,f) our captured sequences .
@@ 5.1 Comparison with the State of the Art
We evaluate our method on the public datasets, Human3.6M [14], 3DPW [58], EHF [40] and our synthetic dataset. Tab. 1 shows the comparison results of our method with state-of-the-art ones. Note that the methods reported in Tab. 1 include skeleton-based methods and mesh-based methods.
Table 2: The action-wised comparison of different setups on the Human3.6M dataset.
Method GT keypoints CPN Detection
MPJPE P-MPJPE MPJPE P-MPJPE
In-door 48.2 37.9 59.1 45.7
Synth 55.0 40.6 65.8 49.6
Fusion 44.2 34.5 54.6 43.3
Fusion+ND 40.5 32.3 53.9 42.4
Although the skeleton-based methods achieve higher MPJPE/P-MPJPE scores, they are limited in the ability to capture vertex-to-ground contact information and human kinematics. Our algorithm yields comparable results to recent state-of-the-art mesh-based methods and presents several advantages such as real-time efficiency, full-body motion capture capability and plausible ground-contact constraints. As shown in Fig. 5, we demonstrate that our method generalizes well in three public datasets and our captured sequences.
@@ 5.2 Ablation Studies
In this section, we conduct ablation studies using various settings to validate each component of our proposed method. To validate the effectiveness of the proposed synthetic dataset, we adopt our network without the neural motion descent module as the Baseline method. As reported in Tab. 2, we conduct ablation experiments with four distinct settings: “In-door”: training the Baseline method on Human3.6M only; “Synth”: training the Baseline method on the proposed synthetic datasets only; “Fusion”:training the Baseline method on both the Human3.6M and synthetic datasets; “Fusion+ND”: training our completed solution with the neural motion descent module on both datasets. The fusion training schedule demonstrates superior accuracy compared to training solely on the Human3.6M or synthetic dataset, as observed in both CPN-detected 2D inputs and ground truth (GT) 2D inputs. The reason is that training solely on Human3.6M leads to overfitting due to the limited motion diversity, while training solely on the synthetic dataset may result in a domain gap between the pseudo motion pair and dataset annotations. Furthermore, we present the results of our complete pipeline with the neural descent module, which shows improved performance.
Figure 6: The ablation study of foot floating (FF) and ground penetration (GP). We vary the threshold from 1cm to 3cm to calculate the corresponding FF and GP metrics.
@@
Table 3: The ablation study on our synthetic dataset.
Temp Desc Body Hand Global FF GP
42.3 11.6 60.0 6.69 6.71
36.2 10.3 54.8 8.3 4.1
39.2 9.6 52.3 0.081 0.23
36.1 9.1 47.4 0.058 0.086

### v3

@@ 4.2 Sequential Full-body Motion Initialization
For better body-hand compatibility, we exploit the cross-attention mechanism to facilitate the motion context sharing during the body and hand recovery. Specifically, we first obtain the initial body features $F_{body}$ and hand features $F_{hand}$ from the temporal encoders and map them as Query, Key, Value matrices in forms of $Q_{body/hand}$, $K_{body/hand}$, and $V_{body/hand}$, respectively. Then we update the body features $F^{\prime}_{body}$ and hand features $F’_{hand}$ as follows:
$F^{\prime}_{body}$ $=V_{body}+Softmax(\frac{Q_{hand}K_{body}^{\top}}{\sqrt{d_{k}}})V_{body},$ (2)
$F^{\prime}_{hand}$ $=V_{hand}+Softmax(\frac{Q_{body}K_{hand}^{\top}}{\sqrt{d_{k}}})V_{hand}.$
@@
The updated features $\{F^{\prime}_{body},F^{\prime}_{hand}\}$ can be further utilized in the contact indicators $ind$ and serve as the temporal context in the Neural Descent module, as will be described shortly. In our experiments, we demonstrate that the feature fusion in Eq. 2 can make two tasks benefit each other to produce more comparable wrist poses in the full-body model.
@@ 4.3 Contact-aware Neural Motion Descent
At the second stage of our method, the initial motion predictions will be refined to be more accurate and physically plausible with the proposed contact-aware neural motion descent module. As shown in Fig. 2, this module takes the 2D misalignment and body-ground contact status as input and updates motion parameters during iterations.
Misalignment and Contact Calculation.
At the iteration of $i\in\{0,1,...,N\}$, we calculate the 2D misalignment status by projecting the 3D joints on the image plane and calculate the differences between the re-projected 2D joints and the proxy observations: $\mathcal{S}_{proj}=\Pi(J_{i},\{K,R_{H}^{i},T_{H}^{i}\})-\widehat{J}_{2D}$. Here, $\Pi(\cdot)$ denotes the perspective projection function, and $K$ denotes the intrinsic parameter.
@@
For the contact status, we calculate the velocity of 3D joints $v_{xz}^{i}$ in xz-plane and the distance to the ground as $d_{y}^{i}$. Moreover, we also leverage the temporal features from inputs 2D skeletons to predict the contact labels $ind$, which will be used as an indicator to mask the body-ground contact. Then, the contact status of the current predictions can be obtained as $\mathcal{S}_{contact}=ind\odot(v_{xz}^{i},d^{i})$, where $\odot$ denotes the Hadamard product operation.
Motion Update.
After obtaining the contact and misalignment status, we feed them into the neural motion descent module for motion updates. As shown in Fig. 4, the descent module takes the two groups of tensors as input: i) the state group includes the current SMPL parameters in the human coordinate system $\beta^{i},t^{i}_{H},\theta^{i}_{H}$, camera pose $R^{i}_{H},T^{i}_{H}$ and the sequential motion context $F_{seq}=\{F^{\prime}_{body},F^{\prime}_{hand}\}$; ii) the deviation group includes the current misalignment status $\mathcal{S}_{proj}$ and contact status $\mathcal{S}_{contact}$.
@@
A straightforward solution would be using an MLP to process these two groups of tensors. However, the values of these two groups exhibit significant differences. For instance, the values of the state tensors change smoothly while the values of the deviation tensors may change rapidly along with the misalignment and contact status. Simply concatenating them as inputs introduces difficulty in the learning process. Note that the magnitude of the deviation tensors is highly correlated with the parameter updates. When the body model is well-aligned without foot skating or ground penetration, the values of the deviation tensors are almost zero, indicating that the refiner should output zeros to prevent further changes in the pose parameters. Otherwise, the refiner should output larger update values for motion adjustments. To leverage such a correlation property, we exploit a cross-attention module to build a more effective architecture.
As shown in Fig. 4, two fully connected layers are leveraged to process the tensors of the state and deviation groups and produce the Query, Key, and Value for the cross-attention module. In this way, our contact-aware neural motion descent module can effectively learn the relationship between the state and deviation groups and hence produce more accurate motion updates. Moreover, the sequential motion context $F_{seq}$ is also leveraged in our neural descent module to mitigate the depth uncertainty and improve the motion predictions.
@@
Compared with previous work [45, 73, 52, 68], the proposed contact-aware neural motion descent module offers the advantage of freeing us from the heavy cost of explicit gradient calculations or the manual tuning of hyper-parameters during testing. Furthermore, the module is capable of learning human motion priors with contact information from our synthetic dataset, which provides a more suitable descent direction and steps to escape the local minima and achieve faster convergence.
Figure 4: Implementations of the neural descent module.
@@ 4.4 Loss Function
In our solution, the full-body motion recovery module and the contact-aware neural motion descent module are trained sequentially. Benefiting from the proxy-to-motion learning, the ground-truth full-body pose $\theta,g$, and human body shape $\beta_{b}$ can be obtained for supervision from our synthetic dataset. Overall, the objective function of motion recovery can be written as follows:
$\mathcal{L}_{rec}=\mathcal{L}_{3D}+\mathcal{L}_{2D}+\mathcal{L}_{\theta}+\mathcal{L}_{\beta}+\mathcal{L}_{cam}+\mathcal{L}_{consist}+\mathcal{L}_{smooth}$ (3)
Specifically, $\mathcal{L}_{3D}$ involves 3D MPJPE loss and 3D trajectory L1 loss while $\mathcal{L}_{2D}$ is the projected 2D MPJPE loss. $\mathcal{L}_{\theta},\mathcal{L}_{\beta},\mathcal{L}_{cam}$ represents L1 loss between the estimated human pose, shape and camera pose to our synthetic ground truth. $\mathcal{L}_{consist}$ is a L1 loss to restrict the consistency of the local motion outputs $\theta_{H},t_{H}$ of the same 3D motion sequence via different observations from virtual cameras. $\mathcal{L}_{smooth}$ is adopted from  [69] by penalizing the velocity and acceleration between the estimation and the ground truth. For the neural descent module, the objective loss can be written as:
$\begin{cases}\mathcal{L}_{desc}=\sum_{k}{u^{N-k}(\mathcal{L}_{rec}+\mathcal{L}_{contact})}\\ \mathcal{L}_{contact}=\sum_{i}{ind^{gt}\odot(||v_{xz}||_{2}+||d_{y}||_{2})}\\ \mathcal{L}_{ind}=\sum_{i}{Entropy(ind^{gt},ind^{est})}\end{cases}$ (4)
where $k=1,2,...,N$ is the iteration time and $u$ is the decay ratio to emphasize the last iteration. We set $K=3,u=0.8$ in practice. $\mathcal{L}_{contact}$ involves the error of trajectory drifting, foot floating, or ground penetration. $\mathcal{L}_{ind}$ refers to the loss between the predicted contact label to the ground truth.
@@ 5 Experiments
In this Section, we validate the efficacy of our method and demonstrate accurate human motion capture results with physically plausible foot-ground contact in world space.
Dataset. The RICH dataset [11] is collected with a multi-view static camera system and one moving camera that the ground truth 3D human motions can be recovered using spatial stereo geometry. The EgoBody [74] is captured by a multi-camera rig and a head-mounted device, focusing on the social interactions in 3D scenes. Dynamic Human3.6M is a benchmark proposed in  [67] to simulate the moving cameras on Human3.6M [13] by randomly cropping with a small view window around the person.
Metrics. In our experiments, we follow previous work [67] to report various metrics, primarily focusing on the evaluation of the world coordinate system. The WA-MPJPE metric reports the MPJPE after aligning the entire trajectory of both the predicted and GT through Procrustes Alignment. The W-MPJPE metric reports the MPJPE after aligning the first frame of sequences. The PA-MPJPE metric reports the MPJPE error after applying the ground truth trajectories to each frame. The ACCEL metric is used to evaluate the joint acceleration.
Figure 5: Results across different cases in the (a,g) 3DPW [61], (b) EHF [40], and (c) Human3.6M [13] datasets and (d,e,f) internet videos. We demonstrate that our method can recover the accurate and plausible human motions in moving cameras at a real-time performance. Specifically, (g) demonstrates the robustness and the temporal coherence of our method even under the occlusion inputs.
@@ 5.1 Comparison with the State of the Art
[…]
Methods W-MPJPE $\downarrow$ WA-MPJPE $\downarrow$ PA-MPJPE $\downarrow$ ACCEL $\downarrow$
EgoBody dataset
[…]
GLAMR [67] 416.1 239.0 114.3 173.5
Ours 385.5 131.3 73.5 49.6
RICH dataset
[…]
GLAMR [67] 653.7 365.1 79.9 107.7
Ours 629.8 343.6 56.0 25.3
We compare our approach with the state-of-the-art approaches to human motion recovery under dynamic cameras, including GLAMR [67], SLAMHR [63] and PACE [18]. Both the SLAMHR and PACE require a pre-processing SLAM to reconstruct the scene to solve the camera trajectories (refer to Tab. 4). Such a process is time consuming and requires texture-rich backgrounds, which narrows their applications. To validate the effectiveness of the proposed solution, we primarily compare our method with GLAMR, as it also runs without SLAM. We also conduct comparison experiments on the RICH and EgoBody datasets, as shown in Tab. 1. As shown in the table, our method achieves significant improvements over previous solutions in all metrics.
Visual comparisons with previous different solutions are also depicted in Fig. 6 and the video in our supplementary materials, where our method again shows superior results in terms of model-image alignment and foot-ground contact in world space.
Figure 6: Qualitative comparison with previous state-of-the-art methods: (a) PyMAF-X [72], (c) GLAMR [67], (b)(d) Ours.
@@ 5.2 Ablation Studies
We conduct ablation studies to validate the effectiveness of the proposed neural descent method on the Dynamic Human3.6M dataset following the setup of  [67]. As shown in Tab 2, the Neural Descent module can significantly reduce the motion errors in world space.
@@
Table 2: Ablation study of the Neural Descent module on the Dynamic Huaman3.6M dataset.
Neural Descent W-MPJPE $\downarrow$ PA-MPJPE $\downarrow$
w/o 644.8 48.4
w/ 605.4 45.9
We also report the metric of ground penetration [66] (GP) and foot floating (FF) in the Human3.6M [13] dataset. The GP is defined as the percentage of the frames that penetrate to the ground. The FF is defined as the percentage of frames with foot-ground distances far from the given threshold. We report the curves of GP and FF with respect to the distance to ground in Fig. 7 with a logarithmic scale, where we can conclude that the neural descent algorithm can significantly improve the ground contact plausibility.
Figure 7: The ablation study on the percentage of foot floating (FF) and ground penetration (GP). We vary the threshold from 1cm to 3cm to calculate the corresponding FF and GP metrics.
5.3 Runtime
It is also worth noting that the proposed method has improved the speed by an order of magnitude compared to the previous methods. The speeds of different methods are reported in Tab. 4. Our method can reach real-time performance at 30 FPS in a laptop with RTX4060, which is very promising to enable various applications related to virtual humans.
Table 3: Runtime comparison with the state-of-the-art methods.
Method SLAMHR PACE GLAMR Ours
FPS 0.04 2.1 2.4 30
6 Conclusion
In this paper, we present ProxyCap, a real-time monocular full-body motion capture approach with physically plausible foot-ground contact in world space. We leverage a proxy dataset based on 2D skeleton sequences with accurate 3D rotational motions in world space. Based on the proxy data, our network learns motions from a human-centric perceptive, which enhances the consistency of human motion predictions under different camera trajectories. For more accurate and physically plausible motion capture, we further propose a contact-aware neural motion descent module so that our network can be aware of foot-ground contact and motion misalignment. Based on the proposed solution, we demonstrate a real-time monocular full-body capture system under moving cameras.
Limitations.
As our method recovers 3D motion from 2D joint observations, the depth ambiguity issue remains especially when the person is captured in nearly static poses.
Supplementary Material
7 Implementation Details
7.1 Human-to-World Coordinate Transformation
We estimate the local human pose of each frame in our proxy-to-motion network, then we transform it into the global world space. In the first frame, the accumulated translation of human in x-z plane $t_{xz}$ is set to zero. For the later human-space estimations, we firstly rotate the front axis of camera in human space to align the y-z plane by $R_{front}$. Denote the target parameters of camera in world space as $R_{W},T_{W}$ and the predicted parameters in human space as $R_{H},T_{H}$. We have $R_{W}=R_{H}\cdot R_{front}$, $T_{W}=-R_{W}\cdot(R_{front}^{T}\cdot(-R_{H}^{T}\cdot T_{H})+t_{xz})$. And the human orientation should also be rotate in the same time to maintain relative stillness: $\theta_{W}(root)=R_{front}^{T}\cdot\theta_{H}(root)$. The world translation can be calculated by $t_{W}=R_{front}^{T}\cdot(t_{H}+J_{root})-J_{root}+t_{xz}$. Here $J_{root}$ is the root joint of SMPL model in T-pose to eliminate the error resulted from the root-misalignment with the original point of SMPL model. Finally, the accumulated translation of human in x-z plane $t_{xz}$ is updated by $t_{xz}^{t+1}=t_{xz}^{t}+R_{front}^{T}\cdot t_{H}$.
7.2 Partial Dilated Convolution
Figure 8: Illustration of human recovery. The left part (a) depicts the original dilated convolution backbone of  [41], while the right part (b) illustrates our proposed partial dilated convolution architecture. Our approach selectively excludes corrupted data from input signals, allowing us to extract motion features more accurately and effectively. Specifically, detection failure (denoted as green mosaic squares) may occur during extremely fast motion or severe occlusion situations, while our architecture will cut off the connection from corrupted data to prevent disturbance transfer during network forward processing.
It should be noted that the hand area, being an extremity, is more prone to being affected by heavy motion blur and severe occlusions, resulting in missing detections. Simply setting the corrupted data to zero is not a viable solution as the original convolution kernel is unable to distinguish between normal data and corrupted data, leading to a significant reduction in performance as noise is propagated through the network layers.
To overcome this challenge, we employ partial convolution [25] to enhance our 1D dilated convolution framework. As illustrated in Fig. 8, rather than indiscriminately processing the input signals as in the original convolution operator, we utilize a mask-weighted partial convolution to selectively exclude corrupted data from the inputs. This enhances the robustness of hand recovery in scenarios involving fast movement and heavy occlusion. Specifically, the latent code $X_{0}$ is initially set as the concatenation of the $(x,y)$ coordinates of the $J$ joints for each frame $f\in 1,2,...,L$, while the mask $M_{0}$ is initialized as a binary variable with values of 0 or 1, corresponding to the detection confidence. Then we integrated the 2D partial convolution operation and mask update function from  [25] into our 1D dilated convolution framework:
$\begin{cases}M_{k+1}=I\left(sum(M_{k})>0\right)\\ X_{k+1}=M_{k+1}\left(W_{k}^{\top}\left(X_{k}\odot M_{k}\right)\frac{size(M_{k})}{sum(M_{k})}+b_{k}\right)\end{cases}$ (5)
where $W$ and $b$ denotes the convolution filter weights and bias at layer $k$, and $\odot$ denotes the element-wise multiplication. Furthermore, in the training procedure, half of the sequential inputs are randomly masked to simulate detection failures that may occur in a deployment environment.
7.3 Training
We train our network using the Adam optimizer with a learning rate of 1e-4 and a batch size of 1024 in NVIDIA RTX 4090. We adopt a two-stage training scheme. Firstly, we train our proxy-to-motion initialization network (Sec. 4.2) for 50 epochs. Subsequently, we fix the weights of the motion recovery network and train the neural descent network (Sec. 4.3) for another 50 epochs.
7.4 Proxy Dataset
We conduct the training process on our synthetic proxy dataset. The 3D body rotational motions of the training set are sampled from AMASS [29]: [ACCAD, BMLmovi, BMLrub, CMU, CNRS, DFaust, EKUT, Eyes Japan Dataset, GRAB, HDM05, HumanEva, KIT, MoSh, PosePrior, SFU, SOMA, TCDHands, TotalCapture, WEIZMANN] and [SSM, Transitions] are for testing. Otherwise, for the generation of hand motions, we adopt the same dataset division of InterHand [33]. Then We animate the SMPL-X mesh and generate virtual cameras to obtain the pseudo 2D labels.
8 Computational Complexity
In this section, we compare the inference speed of our method. Our real-time monocular full-body capture system can be implemented on a single Laptop (NVIDIA RTX 4060 GPU). Specifically, for the 2D pose estimator, we leverage the off-the-shelf Mediapipe [26] and MMPose and re-implement it on the NVIDIA TensorRT platform. We report the inference time of each module in Table. 4.